A) On Earth (a = 10 m/s/s) you throw a ball downward from the top of a 75 m cliff. The ball takes 3 seconds to strike the ground. How fast was the ball going when you released it?

B) A sports car goes racing down the street and skids to a stop. You measure the length of the skid to be 50 m. This car's brakes give it a 4 m/s/s stopping acceleration. How fast was the car going before braking?

C) You roll a ball up a hill, releasing the ball at 12 m/s. The angle of the hill gives the ball an acceleration of 6.0 m/s/s, directed down the hill. How long will it take the ball to return to you at the bottom of the hill?
Answer #1

A) find initial velocity

time =3s and height =75m

Use Formula s=ut+1/2at^{2}

S1/2gt2

75m = u * 3s+ 0.5 * 10m/s2(3s)2

75 -3u, +0.5 10 32

75=3u_{y}+45

30=3u_{y}

ANSWER: {color{Red} u_{y}=10m/s}

======================

B)distance covered while braking =50m

deceleration =4m/s2

final velocity =0m/s

Use Formula v^{2}-u^{2}=2as

(0m/s)^{2}-u^{2}=-2*4m/s^{2}*50m

0-u^{2}=-2*4*50

u=sqrt{2*4*50}

ANSWER: {color{Red} u=20m/s}

===================

C)Use formula

v=u+at

Rolling up

Final velocity is zero

so 0m/s=12m/s-6m/s^{2}*t_{1}

0=12-6*t_{1}

t_{1}=2s

Use formula v^{2}-u^{2}=2as to find distance covered

2 6m/s * s

0-12^{2}=-2*6*s

12^{2}=12s

s=12m

--------------

Rolling down

Use formula s=ut+1/2at^{2}

initial velocity is zero

12m=0+1/2*6m/s^{2}*t_{2}^{2}

12=1/2*6*t_{2}^{2}

12=3t_{2}^{2}

4ー

t_{2}=2s

========

total time =t1 +t2 =2s +2s

ANSWER: t = 4s

====================